OpenStudy (anonymous):

How do I justify my steps? 24 - x = -16

OpenStudy (anonymous):

@CANDESTIEL 24 - x = -16 -x+24=-16 -x + 24 - 24 = -16 - 24 -x = -40 -x/-1 = -40/-1 x = 40

OpenStudy (anonymous):

Alrighty, we have 24 - x = -16 24 - x = -16 +16 +16 ________________ 40 - x = 0 +x +x ______________ 40 = x Justification: You add 16 to both sides in order to get the numbers together and isolate the x. After this, we add x to both sides to isolate the variable. the justification for all of this is merely to solve for x. Hope I helped.